OpenStudy (anonymous):

Help with math! Medal and fan <3

OpenStudy (anonymous):

A pair of parallel lines is cut by a transversal, as shown below:

OpenStudy (anonymous):

What is the measure of angle x? 60 degrees 70 degrees 80 degrees 100 degrees

OpenStudy (anonymous):

@AlexandervonHumboldt2 @luckycoins888 Please help!!!!!!

OpenStudy (alexandervonhumboldt2):

x+20=80 as they are lternate interior angles

OpenStudy (alexandervonhumboldt2):

x=80-20

OpenStudy (anonymous):

X = 60?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

Can you help me with a few more?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

Which relationship is always correct for the angles x, y, and z of triangle ABC? x + z = y y + z = x x + y + z = 180 degrees x + y + z = 90 degrees

OpenStudy (alexandervonhumboldt2):

y + z = x i tihnk

OpenStudy (anonymous):

Do you have an explanation? ^

OpenStudy (alexandervonhumboldt2):

there is a theorem that states that two angles in sum are equal to the 180-third angle. x=180-third anlge so y+z=x

OpenStudy (anonymous):

Ok! Next one

OpenStudy (anonymous):

What is the measure of angle CED? p over 2, because Triangle ABC is congruent to triangle DCE p + 70 degrees, because Triangle ABC is similar to triangle DCE 110 degrees − p, because Triangle ABC is congruent to triangle DCE p, because Triangle ABC is similar to triangle EDC

OpenStudy (anonymous):

Sorry it got blocked out one second 110 degrees - p because

OpenStudy (alexandervonhumboldt2):

p, because Triangle ABC is similar to triangle EDC We have 2 given angles that are congruent so we can state that hey are similar

OpenStudy (anonymous):

A pair of parallel lines is cut by a transversal, as shown below: Which of the following best represents the relationship between angles x and y? x = y x = 180 degrees − y y = 180 degrees − x x = 2y

OpenStudy (alexandervonhumboldt2):

they are alternate interoior angles alternate interoior angle are congruent

OpenStudy (anonymous):

So the answer would be a? X = Y?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

Ok! Last one!

OpenStudy (anonymous):

The figure shows two parallel lines KL and NO cut by the transversals KO and LN. Which statement best explains the relationship between Triangle KLM and Triangle ONM? Triangle KLM is similar to triangle ONM because measure of angle 3 equals measure of angle 4 and measure of angle 1 equals measure of angle 5 Triangle KLM is similar to triangle ONM because measure of angle 3 equals measure of angle 6 and measure of angle 1 equals measure of angle 4 Triangle KLM is congruent to triangle ONM because measure of angle 3 equals measure of angle 4 and measure of angle 1 equals measure of angle 5 Triangle KLM is congruent to triangle ONM because measure of angle 3 equals measure of angle 6 and measure of angle 1 equals measure of angle 4

OpenStudy (alexandervonhumboldt2):

Triangle KLM is similar to triangle ONM because measure of angle 3 equals measure of angle 4 and measure of angle 1 equals measure of angle 5

OpenStudy (anonymous):

Can you explain? @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

3=4 becasuee they are vertical 1=5 becasue thay are alternate interior angles we have enought inmormation to state that triangle are similar

OpenStudy (anonymous):

Thanks I got 100%!